# OpenTCM: A GraphRAG-Empowered LLM-based System for Traditional Chinese Medicine Knowledge Retrieval and Diagnosis

### v2

@@ Abstract
Traditional Chinese Medicine (TCM) represents a rich repository of ancient medical knowledge that continues to play an important role in modern healthcare. Due to the complexity and breadth of the TCM literature, the integration of AI technologies is critical for its modernization and broader accessibility. However, this integration poses considerable challenges, including the interpretation of obscure classical Chinese texts and the modeling of intricate semantic relationships among TCM concepts. In this paper, we develop OpenTCM, an LLM-based system that combines a domain-specific TCM knowledge graph and Graph-based Retrieval-Augmented Generation (GraphRAG). First, we extract more than 3.73 million classical Chinese characters from 68 gynecological books in the Chinese Medical Classics Database, with the help of TCM and gynecology experts. Second, we construct a comprehensive multi-relational knowledge graph comprising more than 48,000 entities and 152,000 interrelationships, using customized prompts and Chinese-oriented LLMs such as DeepSeek and Kimi to ensure high-fidelity semantic understanding. Last, we integrate OpenTCM with this knowledge graph, enabling high-fidelity ingredient knowledge retrieval and diagnostic question-answering without model fine-tuning. Experimental evaluations demonstrate that our prompt design and model selection significantly improve knowledge graph quality, achieving a precision of 98. 55% and an F1 score of 99. 55%. In addition, OpenTCM achieves mean expert scores of 4.5 in ingredient information retrieval and 3.8 in diagnostic question-answering tasks, outperforming state-of-the-art solutions in real-world TCM use cases.
@@ I Introduction
Traditional Chinese Medicine (TCM) represents a rich repository of ancient medical knowledge that continues to hold significant value in modern healthcare systems [1]. Due to the complexity of the literature and the practice of TCM, many researchers have tried to build systems to assist with consultation and diagnosis for TCM physicians. Current TCM knowledge retrieval and diagnosis systems predominantly often follow a paradigm of fine-tuning general-purpose Large Language Models (LLMs), such as ChatGPT [2] or Claude [3], on domain-specific datasets. However, they often face significant challenges such as the complexity of parsing classical Chinese texts, the intricate relationships among TCM concepts (such as ingredients, symptoms, and diagnostic principles), and the absence of structured knowledge such as knowledge graphs and retrieval-augmented generation (RAG) [4, 5, 6]. In particular, ancient TCM literature is often written in archaic language, making them difficult for modern practitioners and computational models to interpret. Furthermore, existing approaches primarily focus on modern medical literature, overlooking the foundational corpus of classical TCM. Although recent advances in LLMs have proven to be effective in medical knowledge democratization, the direct application of general-purpose LLMs to TCM still faces significant limitations, such as hallucination in prescription generation and semantic understanding of classical Chinese.
@@
To address these limitations, we propose OpenTCM, an LLM-based TCM knowledge retrieval and diagnosis system that integrates domain-specific knowledge with Retrieval-Augmented Generation with Graphs (GraphRAG) [7]. First, we extract over 3.73 million ancient Chinese characters from 68 gynecological books sourced from the Chinese Medical Classics Database [8], with verification from experienced TCM and gynecology experts. Next, we construct a comprehensive multi-relational knowledge graph containing over 48,000 entities and 152,000 interrelationships. This graph spans key TCM concepts, including 3,700+ ingredients, 14,000+ diseases, 17,000+ symptoms and treatments, and 65,000+ ingredient references. The construction process leverages customized prompts and domain-adapted LLMs, such as DeepSeek [9] and Kimi [10], to ensure accurate semantic understanding of classical Chinese medical literature. Compared to general-purpose prompts and LLMs, our approach achieves significant improvements in knowledge graph quality, with a precision of 98.55%, recall of 99.60%, F1-score of 99.55%, and accuracy of 98.17%. In addition, our experiments show that OpenTCM achieves a Mean Expert Score (MES), accuracy, and inter-rater agreement of 4.5, 89.6%, and 0.81 for ingredient retrieval, and 3.8, 75.1%, 0.78 for diagnostic question-answering, respectively. By eliminating the need for model fine-tuning, OpenTCM leverages retrieval-augmented generation to provide accurate, context-aware responses with minimal computational overhead. Our structured knowledge graph also mitigates hallucination risks and enhances reasoning over complex TCM relationships, offering a scalable and efficient alternative to traditional methods.
@@
  • We integrate the knowledge graph with GraphRAG to enhance reasoning capabilities of OpenTCM over complex TCM interrelationships while maintaining computational efficiency.
  • Our experimental evaluation demonstrates that our system significantly outperforms existing solutions, achieving a precision of 98.55% in knowledge graph construction task, an MES of 4.5 and accuracy of 89.6% in ingredient retrieval, and an MES of 3.8 and accuracy of 75.1% in diagnostic question-answering.
@@ III System Overview
The process starts with collecting gynecological books from the Chinese Medical Classics database [8]. These medical literature, covering obstetrics, gynecology, and fertility, are carefully reviewed and corrected by TCM and gynecology medical experts to ensure data quality. Next, a multi-relational knowledge graph is constructed using well-customized prompts and advanced domain-adapted LLMs like Deepseek and Kimi. This knowledge graph captures key TCM elements such as ingredients, symptoms, and treatments, forming a structured representation of TCM. Last, the constructed knowledge graph serves as the backbone for FastGPT, enabling the implementation of GraphRAG. This integration allows OpenTCM to offer powerful features such as TCM ingredient search and diagnostic question-answering without requiring pre-training or fine-tuning of the LLMs, significantly reducing computational overhead while maintaining high performance.
@@ III-C Graph-based Retrieval-Augmented Generation
We use GraphRAG [7, 20, 21] to integrate graph-structured data into OpenTCM to enhance retrieval and generation processes. While conventional RAG approaches predominantly processes unstructured text or images, GraphRAG leverages the rich structural relationships embedded within a knowledge graph to improve both information access and reasoning capabilities. This is particularly valuable for TCM, which features complex inter-dependencies among ingredients, symptoms, treatments, and syndromes.
@@
These adaptations enhances our system’s performance by relying on a domain-specific graph structure. Therefore, OpenTCM mitigates hallucination and noise commonly observed in generic LLM outputs. This design supports high-precision tasks such as ingredient knowledge retrieval (accuracy: 89.6%) and diagnostic question-answering (accuracy: 75.1%), while reducing the need for model fine-tuning. The integration of GraphRAG with the TCM knowledge graph thus forms the core of OpenTCM’s ability to deliver accurate, interpretable, and context-aware responses in real-world medical scenarios.
@@ IV-A1 Knowledge Graph Construction Baselines
  • General-Purpose LLMs: We utilize ChatGPT-4o [2] and Claude-2 [22] as representatives of general-purpose large language models.
@@ IV-A2 Real-World Application Baselines
  • General-Purpose LLMs: ChatGPT-4o [2] and Claude-2 [22] are used to assess the performance of general-purpose models in TCM applications.
  • TCM-Specific LLMs: BianCang [14] and Lingdan [15] are recent models specifically designed for TCM tasks, incorporating domain-specific knowledge and training data.
@@ IV-B Knowledge Graph Construction
Evaluating the quality of the TCM knowledge graph is a cornerstone of this study, given its role as the foundation for OpenTCM’s downstream applications. In the absence of a pre-existing gold standard for TCM knowledge graph construction, we adopt a sampling inspection approach, enlisting 5 domain experts in TCM and Gynecology to assess the generated triples. Our evaluation examines two dimensions: (1) the effectiveness of customized versus general prompts, and (2) the performance of different backbone models, including Chinese-oriented LLMs (i.e., Deepseek [9], Kimi [10]) and more general LLMs (i.e., ChatGPT-4o [2], Claude [22]).
@@ IV-B2 Performance
General Customized
Precision 90.1% 98.55%
Recall 95.9% 99.60%
F1-score 92.3% 99.55%
Accuracy 86.8% 98.17%
@@ IV-C Real-world Applications
Due to the absence of a definitive ground truth, we collect and annotate a dataset consisting of 200 ingredient information search queries and 200 diagnostic consultation questions. Each query is processed by OpenTCM and several baseline systems, including general-purpose LLMs (GPT-4, Claude-2) and recently proposed TCM-specific LLMs (BianCang, Lingdan). The responses are scored by five medical experts on a 5-point Likert scale (1 = irrelevant, 5 = highly relevant).
IV-C1 Metric
We use three metrics for evaluation:
  • Mean Expert Score (MES): The average score assigned by domain experts to evaluate the relevance and correctness of the retrieved or generated information.
  • Accuracy: The proportion of expert-labeled correct responses over the total number of queries.
  • Inter-Rater Agreement (IRA): A measure of consistency among expert evaluators to ensure reliability of the results.
@@ IV-C2 Ingredient Information Retrieval
This task evaluates the model’s ability to return accurate and comprehensive information about ingredients, such as therapeutic properties and relationships to treatments. Queries range from basic lookups (e.g., “What is the function of Dang Gui?”) to more complex relational queries (e.g., “Which ingredients treat irregular menstruation alongside abdominal pain?”).
@@
GPT-4 4.0 84.3% 0.78
Claude-2 3.9 83.1% 0.76
Lingdan 4.3 87.2% 0.79
BianCang 4.1 85.5% 0.78
OpenTCM 4.5 89.6% 0.81
@@ IV-C3 Diagnostic Question Answering
This task assesses the system’s diagnostic reasoning capabilities in clinical consultation scenarios (e.g., “What medication should be used for postpartum headaches?” or “How to treat Qi deficiency with dampness?”). An example is illustrated in Figure 4.
Figure 4: An example case of OpenTCM diagnostic question-answering.
@@
GPT-4 3.4 69.0% 0.74
Claude-2 3.5 70.2% 0.75
Lingdan 3.7 74.5% 0.79
BianCang 3.6 72.4% 0.75
OpenTCM 3.8 75.1% 0.78
@@ VI Conclusion
By integrating knowledge graph and GraphRAG, we developed OpenTCM, a LLM-based TCM knowledge retrieval and diagnosis system using TCM databases that consists of 68 TCM books with 6700+ chapters and 3,7+ million ancient Chinese characters. Experimental results demonstrate that OpenTCM achieves a precision of 98.55% in knowledge graph construction, an MES of 4.5 and accuracy of 89.6% in ingredient retrieval, and an MES of 3.8 and accuracy of 75.1% in diagnostic question-answering, outperforming existing solutions.

### v3

@@ Abstract
Traditional Chinese Medicine (TCM) represents a rich repository of ancient medical knowledge that continues to play an important role in modern healthcare. Due to the complexity and breadth of the TCM literature, the integration of AI technologies is critical for its modernization and broader accessibility. However, this integration poses considerable challenges, including the interpretation of obscure classical Chinese texts and the modeling of intricate semantic relationships among TCM concepts. In this paper, we develop OpenTCM, an LLM-based system that combines a domain-specific TCM knowledge graph and Graph-based Retrieval-Augmented Generation (GraphRAG). First, we extract more than 3.73 million classical Chinese characters from 68 gynecological books in the Chinese Medical Classics Database, with the help of TCM and gynecology experts. Second, we construct a comprehensive multi-relational knowledge graph comprising more than 48,000 entities and 152,000 interrelationships, using customized prompts and Chinese-oriented LLMs such as DeepSeek and Kimi to ensure high-fidelity semantic understanding. Last, we integrate OpenTCM with this knowledge graph, enabling high-fidelity ingredient knowledge retrieval and diagnostic question-answering without model fine-tuning. Experimental evaluations demonstrate that OpenTCM achieves mean expert scores (MES) of 4.378 in ingredient information retrieval and 4.045 in diagnostic question-answering tasks, outperforming state-of-the-art solutions in real-world TCM use cases.
@@ I Introduction
Traditional Chinese Medicine (TCM) represents a rich repository of ancient medical knowledge that continues to hold significant value in modern healthcare systems [1]. Due to the complexity of the literature and the practice of TCM, many researchers have tried to build systems to assist with consultation and diagnosis for TCM physicians. Current TCM knowledge retrieval and diagnosis systems predominantly often follow a paradigm of fine-tuning general-purpose Large Language Models (LLMs), such as ChatGPT [2] or Claude [3], on domain-specific datasets. However, they often face significant challenges such as the complexity of parsing classical Chinese texts, the intricate relationships among TCM concepts (such as ingredients, symptoms, and diagnostic principles), and the absence of structured knowledge such as knowledge graphs and retrieval-augmented generation (RAG) [4, 5]. In particular, ancient TCM literature is often written in archaic language, making them difficult for modern practitioners and computational models to interpret. Furthermore, existing approaches primarily focus on modern medical literature, overlooking the foundational corpus of classical TCM. Although recent advances in LLMs have proven to be effective in medical knowledge democratization, the direct application of general-purpose LLMs to TCM still faces significant limitations, such as hallucination in prescription generation and semantic understanding of classical Chinese.
@@
To address these limitations, we propose OpenTCM, an LLM-based TCM knowledge retrieval and diagnosis system that integrates domain-specific knowledge with Retrieval-Augmented Generation with Graphs (GraphRAG) [6]. First, we extract over 3.73 million ancient Chinese characters from 68 gynecological books sourced from the Chinese Medical Classics Database [7], with verification from experienced TCM and gynecology experts. Next, we construct a comprehensive multi-relational knowledge graph containing over 48,000 entities and 152,000 interrelationships. This graph spans key TCM concepts, including 3,700+ ingredients, 14,000+ diseases, 17,000+ symptoms and treatments, and 65,000+ ingredient references. The construction process leverages customized prompts and domain-adapted LLMs, such as DeepSeek [8] and Kimi [9], to ensure accurate semantic understanding of classical Chinese medical literature. Compared to general-purpose prompts and LLMs, our approach achieves significant improvements in knowledge graph quality, with a precision of 98.55%, recall of 99.60%, F1-score of 99.55%, and accuracy of 98.17%. In addition, our experiments show that OpenTCM achieves a Mean Expert Score (MES), accuracy, and inter-rater agreement (IRA) of 4.378, 99.0%, and 0.057 for ingredient retrieval, and 4.045, 98.8%, -0.013 for diagnostic question-answering, respectively. By eliminating the need for model fine-tuning, OpenTCM leverages retrieval-augmented generation to provide accurate, context-aware responses with minimal computational overhead. Our structured knowledge graph also mitigates hallucination risks and enhances reasoning over complex TCM relationships, offering a scalable and efficient alternative to traditional methods.
@@
  • We integrate the knowledge graph with GraphRAG to enhance reasoning capabilities over complex TCM interrelationships while maintaining computational efficiency.
  • Our experimental evaluation demonstrates that our system significantly outperforms existing solutions, achieving a precision of 98.55% in knowledge graph construction task, an MES of 4.378 and accuracy of 99.0% in ingredient retrieval, and an MES of 4.045 and accuracy of 98.8% in diagnostic question-answering.
@@ III System Overview
The process starts with collecting gynecological books from the Chinese Medical Classics database [7]. These medical literature, covering obstetrics, gynecology, and fertility, are carefully reviewed and corrected by TCM and gynecology medical experts to ensure data quality. Next, a multi-relational knowledge graph is constructed using well-customized prompts and advanced domain-adapted LLMs like Deepseek and Kimi. This knowledge graph captures key TCM elements such as ingredients, symptoms, and treatments, forming a structured representation of TCM. Last, the constructed knowledge graph serves as the backbone for the database, enabling the implementation of GraphRAG. This integration allows OpenTCM to offer powerful features such as TCM ingredient search and diagnostic question-answering without requiring pre-training or fine-tuning of the LLMs, significantly reducing computational overhead while maintaining high performance.
@@ III-B Knowledge Graph Construction
General Customized
Precision 90.1% 98.55%
Recall 95.9% 99.60%
F1-score 92.3% 99.55%
Accuracy 86.8% 98.17%
@@ III-C Graph-based Retrieval-Augmented Generation
We use GraphRAG [6, 19] to integrate graph-structured data into OpenTCM to enhance retrieval and generation processes. While conventional RAG approaches predominantly processes unstructured text or images, GraphRAG leverages the rich structural relationships embedded within a knowledge graph to improve both information access and reasoning capabilities. This is particularly valuable for TCM, which features complex inter-dependencies among ingredients, symptoms, treatments, and syndromes.
@@
These adaptations enhances our system’s performance by relying on a domain-specific graph structure. Therefore, OpenTCM mitigates hallucination and noise commonly observed in generic LLM outputs. This design supports high-precision tasks such as ingredient knowledge retrieval (accuracy: 99.0%) and diagnostic question-answering (accuracy: 98.8%), while reducing the need for model fine-tuning. The integration of GraphRAG with the TCM knowledge graph thus forms the core of OpenTCM’s ability to deliver accurate, interpretable, and context-aware responses in real-world medical scenarios.
Figure 4: An example case of OpenTCM diagnostic question-answering.
@@ IV-A1 Knowledge Graph Construction Baselines
  • General-Purpose LLMs: We utilize ChatGPT-4 [2] and Claude-2 [20] as representatives of general-purpose large language models.
@@ IV-A2 Real-World Application Baselines
  • General-Purpose LLMs: ChatGPT-4 [2] and Claude-2 [20] are used to assess the performance of general-purpose models in TCM applications.
  • TCM-Specific LLMs: BianCang [13] is a recent LLM specifically designed for TCM tasks, incorporating domain-specific knowledge and training data.
@@ IV-B Knowledge Graph Construction
Evaluating the quality of the TCM knowledge graph is a cornerstone of this study, given its role as the foundation for OpenTCM’s downstream applications. In the absence of a pre-existing gold standard for TCM knowledge graph construction, we adopt a sampling inspection approach, enlisting 4 domain experts in TCM and Gynecology to assess the generated triples. Our evaluation examines two dimensions: (1) the effectiveness of customized versus general prompts, and (2) the performance of different backbone models, including Chinese-oriented LLMs (i.e., Deepseek [8], Kimi [9]) and more general LLMs (i.e., ChatGPT-4 [2], Claude [20]).
@@ IV-C Real-world Applications
Due to the absence of a definitive ground truth, we collect and annotate a dataset consisting of 257 ingredient information search queries and 303 diagnostic consultation questions. Each query is processed by OpenTCM and several baseline systems, including general-purpose LLMs (KIMI) and recently proposed TCM-specific LLMs (BianCang). The system’s responses were evaluated by a panel of four medical experts, comprising two clinicians with more than 20 years of clinical experience and two medical Ph.D. students whose assessments were validated by their respective supervisors. A 5-point Likert scale, ranging from 1 (irrelevant) to 5 (highly relevant), was employed for scoring.
IV-C1 Metrics
We employ three key metrics for evaluating the performance of the system:
  • Mean Expert Score (MES): This metric represents the average score assigned by domain experts to assess the relevance and correctness of the retrieved or generated information. MES quantitatively evaluates the experts’ subjective judgment of the system’s response quality.
  • Accuracy: Accuracy is defined as the proportion of expert-labeled correct responses out of the total number of queries. For the purpose of this evaluation, a response is deemed correct if its expert score is equal to or greater than 3. This metric measures the overall capability of the model to produce acceptable responses.
  • Inter-Rater Agreement (IRA): IRA quantifies the consistency among multiple domain expert evaluators to ensure the reliability of the assessment results. For the calculation of IRA, expert ratings are binarized: scores greater than or equal to 3 are categorized as ”correct” (or ”agreement”), while scores less than 3 are categorized as ”incorrect” (or ”disagreement”). IRA reflects the objectivity of the evaluation by measuring the extent to which evaluators agree on these binary classifications.
@@ IV-C2 Ingredient Information Retrieval
This task evaluates the model’s ability to return accurate and comprehensive information about ingredients such as “Danggui” and “Renshen”.
@@
Figure 5: Score distribution of different models for Ingredient Information Retrieval across four experts. OpenTCM shows significantly higher concentration at Score 4 and 5.
@@
KIMI 2.691 53.9% -0.154
BianCang 2.616 47.1% -0.166
OpenTCM 4.378 99.0% 0.057
@@ IV-C3 Diagnostic Question Answering
This task assesses the system’s diagnostic reasoning capabilities in clinical consultation scenarios (e.g., “How to treat menstrual disorders?” or “How to treat Qi deficiency with dampness?”). An example is illustrated in Figure 4.
@@
KIMI 3.043 50.2% -0.331
BianCang 2.455 44.7% -0.275
OpenTCM 4.045 98.8% -0.013
Figure 6: Score distribution of different models for Diagnostic Question Answering across four experts. OpenTCM responses receive notably higher expert ratings.
@@ VI Conclusion
OpenTCM represents a significant advancement in applying LLMs to Traditional Chinese Medicine. By combining a meticulously constructed knowledge graph with GraphRAG, it achieves high accuracy and efficiency in knowledge retrieval and diagnostic tasks, paving the way for broader adoption of TCM practices.